### v1

# On planar sections of the dodecahedron

Abstract.
In the analysis of three-dimensional biological microstructures such as organoids, microscopy frequently yields two-dimensional optical sections without access to their orientation. Motivated by the question of whether such random planar sections determine the underlying three-dimensional structure, we investigate a discrete analogue in which the ambient structure is the vertex set of a Platonic solid and the observed data are congruence classes of planar intersections. For the regular dodecahedron with vertex set $V$, we define the planar statistic of a subset $X\subseteq V$ of vertices as the distribution of isometry types of inclusions $\Pi\cap X\subseteq V$, and ask whether this statistic determines $X$ up to planar congruence. We show that this is not the case: there exist two non-congruent $7$-element subsets with identical planar statistics.
As a consequence, there exists two polytopes in $\mathbb{R}^{3}$, whose distribution of isometry classes of two-dimensional intersections is identical, while the polytopes are not themselves isometric. This result is an analogue of classical non-uniqueness phenomena in geometric tomography.
1. Introduction
High-resolution imaging techniques in biological systems frequently produce planar sections of inherently three-dimensional structures. A notable example appears in organoid research, where multicellular aggregates are cultivated in three-dimensional media, yet standard microscopy visualizes them in two-dimensional optical slices, [2, 3, 6]. When we think of one two-dimensional slice through a three-dimensional space filled with more or less identical objects each in a random position and orientation, then we effectively obtain a distribution of random slices of one fixed object without information on what slice exactly we are looking at. More precisely, the resulting observed data consists of a distribution of planar point configurations or shapes up to planar congruence. A fundamental question, both theoretically and for interpretation of imaging data, is whether this distribution encodes the underlying three-dimensional structure uniquely. The problem is rather different from problems in $X$-ray image analysis or geometric tomography, where the position of the intersecting plane is part of the data that can be analysed.
The aim of this work is to study this question in a highly controlled discrete setting. We fix a finite vertex set $V$ of a Platonic solid in $\mathbb{R}^{3}$ and consider subsets $X\subset V$ as discrete models for three-dimensional structures. Planes intersecting $V$ produce subset inclusions $\Pi\cap X\subseteq V$ of planar sets. Only the Euclidean congruence types of these planar configurations are assumed observable, and each plane is unlabelled. Thus the observable information associated to $X$ is merely the distribution of planar congruence classes of all such intersections.
We refer to this distribution as the planar statistic of $X$, see Definition 1. Our primary question is whether two subsets with identical planar statistics must be congruent in $\mathbb{R}^{3}$. By analogy to results of Mallows and Clark [4] on intercept distributions in the plane, non-uniqueness is expected. We establish such non-uniqueness in the case of the regular dodecahedron, see Theorem 3.
This also implies that there exist two three-dimensional convex bodies whose distribution of isometry classes of two-dimensional intersections is identical, while the two convex bodies are not themselves isometric, see Theorem 4.
Let us explain our approach and the relation to previous work in more detail. Our strategy is closely analogous in spirit to the classical example of Mallows and Clark [4, 5], who exhibited two convex dodecagons in the plane with identical chord–length distributions but which are not congruent. Their construction naturally splits into a combinatorial step and a geometric encoding step.
In the combinatorial step, one starts from the regular octagon and considers its eight sides or, equivalently, the eight isosceles sectors with apex at the centre. Mallows and Clark choose two subsets $X$ and $Y$ of four sides each with the property that the multiset of unordered pairs of sides in $X$ is, up to Euclidean isometry of the octagon, the same as the multiset of unordered pairs of sides in $Y$. In modern language, $X$ and $Y$ form a non-congruent homometric pair in the cyclic graph $C_{8}$ that records side adjacency. At this level the construction is purely discrete: one arranges that the “incidence statistics” of $X$ and $Y$ inside the octagon agree, even though the subsets themselves are not related by a symmetry.
The geometric step then promotes this combinatorial configuration to a pair of convex subsets of the plane. To each side in $X$, respectively in $Y$, one attaches a small congruent isosceles triangular hat pointing outward, thereby turning the octagon into two convex dodecagons $K_{X}$ and $K_{Y}$. The caps are chosen so small that every chord of $K_{X}$ or $K_{Y}$ either coincides with a chord of the original octagon, or differs from it only in a tiny segment inside one of the caps. The contribution of a cap to the chord–length distribution depends only on which side it sits on, so the matching of unordered pairs of sides in $X$ and $Y$ forces the chord–length distributions of $K_{X}$ and $K_{Y}$ to coincide, although the polygons themselves are not congruent.
The present dodecahedral construction follows the same two-step pattern in a three-dimensional setting. Here the regular octagon is replaced by the regular dodecahedron $D$ with vertex set $V$, and the role of chords is taken over by planar sections through vertex–planes. The first, combinatorial step is the search for two subsets $S,T\subset V$ with identical planar statistics in the sense of Definition 1. Equivalently, for every vertex–plane $\Pi$ the inclusion $\Pi\cap S\subset V$ has, up to isometry of $D$, the same type and multiplicity distribution as $\Pi\cap T\subset V$. Theorem 3 shows that such a non-congruent pair $(S,T)$ exists inside the vertex–plane incidence structure of the dodecahedron.
In the second step we again encode this discrete data into convex bodies by a local modification along selected faces. Starting from $D$, we cut off a small congruent triangular pyramid at each vertex $v\in X$, obtaining a polytope $K_{X}$ for $X\subseteq V$. When an affine plane $H$ meets $D$ near some vertices, the local shape of the section $K_{X}\cap H$ in a neighbourhood of those vertices is completely determined, up to isometry, by the configuration $\Pi\cap X$, where $\Pi$ is the nearby vertex–plane through the intersected vertices. The distribution of isometry classes of planar sections of $K_{X}$ can therefore be written as a convex combination of contributions coming from the finitely many planar intersection types $\Pi\cap X\subset V$, with coefficients prescribed precisely by the planar statistic $\mathrm{PS}(X)$.
From this viewpoint, Theorem 4 is the exact three-dimensional analogue of the Mallows–Clark theorem: the equality $\mathrm{PS}(S)=\mathrm{PS}(T)$ ensures that the section distributions of $K_{S}$ and $K_{T}$ coincide, while the underlying subsets $S$ and $T$ and hence the patterns of truncated vertices of $D$ are not related by any isometry.
Beyond the discrete-geometric questions addressed here, our section statistics interface naturally with classical inverse problems in geometric tomography and integral geometry [1, 7], where distributions of slices, chords, and projections play a central role. Related stereological ideas go back at least to Wicksell’s corpuscle problem [8] and remain fundamental in modern microscopy when only section data are available.
2. Main results and model
Let $V$ be the 20 vertices of a regular dodecahedron in the golden-ratio model. Let’s write that out in full detail for convenience of the reader. Here, we set $\varphi=\frac{1+\sqrt{5}}{2}.$
idx $(x,y,z)$ idx $(x,y,z)$ idx $(x,y,z)$ idx $(x,y,z)$
0 (1,1,1)(-1,-1,-1) 1 (1,1,1)(-1,-1,1) 2 (1,1,1)(-1,1,-1) 3 (1,1,1)(-1,1,1)
4 (1,1,1)(1,-1,-1) 5 (1,1,1)(1,-1,1) 6 (1,1,1)(1,1,-1) 7 $(1,1,1)$
8 $(0,\varphi^{-1},\varphi)$ 9 $(0,\varphi^{-1},-\varphi)$ 10 $(0,-\varphi^{-1},\varphi)$ 11 $(0,-\varphi^{-1},-\varphi)$
12 $(\varphi^{-1},\varphi,0)$ 13 $(-\varphi^{-1},\varphi,0)$ 14 $(\varphi^{-1},-\varphi,0)$ 15 $(-\varphi^{-1},-\varphi,0)$
16 $(\varphi,0,\varphi^{-1})$ 17 $(\varphi,0,-\varphi^{-1})$ 18 $(-\varphi,0,\varphi^{-1})$ 19 $(-\varphi,0,-\varphi^{-1})$
We depict those vertices in a Schlegel diagram as usual. The full set is depicted as follows:
Figure 1. Schlegel diagram with all 20 vertices labeled.
For $X\subset V$ the planar statistic $\mathrm{PS}(X)$ is the multiset, over all planes $\Pi$ with $|\Pi\cap V|\geq 3$, of planar congruence classes of subset inclusions $\Pi\cap X\subseteq V$, seen as a point sets in $\mathbb{R}^{3}$. More precisely, let $D\subset\mathbb{R}^{3}$ be the dodecahedron as above with vertex set $V$. A vertex-plane is an affine plane $\Pi\subset\mathbb{R}^{3}$ with $|\Pi\cap V|\geq 3$, and we denote by
$\mathcal{H}:=\{\Pi\subset\mathbb{R}^{3}:\Pi\text{ is a vertex-plane of }D\}$
the finite set of all vertex-planes of $D$.
For a subset $X\subset V$ and a vertex-plane $\Pi\in\mathcal{H}$ we consider the inclusion of finite subsets $\Pi\cap X\subseteq V$. Two vertex-planes $\Pi,\Pi^{\prime}\in\mathcal{H}$ are called planarly congruent with respect to $X$ if there exists a Euclidean isometry $g\in\mathrm{Isom}(\mathbb{R}^{3})$ such that $g(V)=V$ and $g(\Pi\cap X)=\Pi^{\prime}\cap X$. We denote the corresponding equivalence relation on $\mathcal{H}$ by $\Pi\sim_{X}\Pi^{\prime}$ and write $\Pi]_{X}$ for the equivalence class of $\Pi$.
Definition 1.
The planar statistic of $X\subset V$ is the multiset $\mathrm{PS}(X)$ of planar congruence classes $\Pi]_{X}$ as $\Pi$ ranges over all vertex-planes in $\mathcal{H}$, counted with multiplicity. In other words, $\mathrm{PS}(X)$ records, for each vertex-plane $\Pi$, the isometry type in $\mathbb{R}^{3}$ of the inclusion $\Pi\cap X\subseteq V$, and forgets the actual position of $\Pi$.
Remark 2.
There is a slight variation of this definition which would record just the isometry type of the inclusion $\Pi\cap X\subseteq\Pi\cap V$ instead of $\Pi\cap X\subseteq V$. This is almost the same amount of information, except that in the case of a square $Q=\Pi\cap V$ with $\Pi\cap X$ two adjacent vertices. In this case, the two distinct embeddings of $Q$ into $V$ give rise to two different planar congruence classes $\Pi]_{X}$.
It is a basic observation that the planar statistic of an image of $X$ under an isometry of $V$, orientation preserving or not, is the same as the one for $X$; i.e. isometric subsets of $V$ yield the same planar statistics. In view of the applications in shape recognition described in the introduction, our initial hope was that the planar statistics $\mathrm{PS}(X)$ allows to recover $X$, but we show that this is not the case. Our main result is the following theorem:
Theorem 3.
There exist two 7-element subsets $S,T\subset V$ that are not congruent in $\mathbb{R}^{3}$, yet $\mathrm{PS}(S)=\mathrm{PS}(T)$. One such pair is $S=\{0,1,2,3,4,11,17\}$ and $T=\{0,1,3,4,5,11,17\}$.
Figure 2. 7-element subsets $S$ (left) and $T$ (right)
In contrast, for all other Platonic solids, in particular for the icosahedron, we find injectivity of planar statistics up to congruence for all sizes of vertex sets – a fact that can be verified computationally. The pair $(S,T)$ is the unique $7$-element pair of subsets up to isometry with the properties as stated in Theorem 3, there are other pairs of size $r$ for $r\in\{8,\dots,13\}.$
Proof.
The proof of Theorem 3 is in part computational: we enumerate all 319 vertex planes and aggregate subconfigurations by planar congruence. This shows that the planar statistics are identical, see below. Since $T=(S\setminus\{2\})\cup\{5\}$, we only need to study the planes in $S$ that contain $2$ and compare to planes in $T$ that contain $5$; this can also be done by direct inspection with a three-dimensional model.
In order to see that $S$ and $T$ are not congruent, observe that each configuration contains a unique path (up to orientation) of length three; this is $(0,11,4,17)$ in both $S$ and $T$. Now observe that the only endpoint of those paths, which does not have other neighbors at distance $\leq 2$ is the vertex $17$ in $S$. Hence, $S$ and $T$ cannot be congruent. ∎
Before we proceed, we need to say a bit more about the possible planar intersections with the vertex set of the dodecahedron. We group the 319 vertex planes by the isometry type of $\Pi\cap V$ and list their frequencies below. The picture shows convex hull of $\Pi\cap V$.
Table 1. List of vertex planes and their frequencies
[Uncaptioned image] (3,0) 60 [Uncaptioned image] (3,0) 20 [Uncaptioned image] (3,0) 60 [Uncaptioned image] (3,0) 60
[Uncaptioned image] (4,0) 30 [Uncaptioned image] (4,0) 30 [Uncaptioned image] (4,0) 15 [Uncaptioned image] (5,0) 12
[Uncaptioned image] (5,0) 12 [Uncaptioned image] (6,0) 20
This list fits with the count of 3-element subsets that we use to generate the planes, indeed we have:
$(20+3\cdot 60)\binom{3}{3}+(15+2\cdot 30)\binom{4}{3}+2\cdot 12\binom{5}{3}+20\binom{6}{3}=1140=\binom{20}{3}$.
Subsequently we display the planar statistics for $S$ and $T$ aggregated by isometry type of the planar embedding $Z\subset V$. The main result follows then simply by comparison of the respective counts. Again, since $T=(S\setminus\{2\})\cup\{5\}$, we only need to count planar configurations in $S$ containing $2$ and compare to planar configurations in $T$ containing $5$. In principle, this can be done by hand. The pictures in the following array show the convex hull of $\Pi\cap V$ with the subset marked. Below the picture one can read $(|\Pi\cap V|,|\Pi\cap S|)$ and the count of the configuration in $S$ and $T$.
Table 2. List of subconfigurations with their frequency
[Uncaptioned image] (3,0) 16 [Uncaptioned image] (3,0) 6 [Uncaptioned image] (3,0) 15 [Uncaptioned image] (3,0) 15
[Uncaptioned image] (3,1) 18 [Uncaptioned image] (3,1) 9 [Uncaptioned image] (3,1) 7 [Uncaptioned image] (3,1) 20
[Uncaptioned image] (3,1) 9 [Uncaptioned image] (3,1) 9 [Uncaptioned image] (3,1) 20 [Uncaptioned image] (3,2) 5
[Uncaptioned image] (3,2) 10 [Uncaptioned image] (3,2) 7 [Uncaptioned image] (3,2) 10 [Uncaptioned image] (3,2) 4
[Uncaptioned image] (3,2) 10 [Uncaptioned image] (3,2) 4 [Uncaptioned image] (3,3) 2 [Uncaptioned image] (3,3) 2
[Uncaptioned image] (3,3) 2 [Uncaptioned image] (4,0) 3 [Uncaptioned image] (4,0) 5 [Uncaptioned image] (4,0) 1
[Uncaptioned image] (4,1) 17 [Uncaptioned image] (4,1) 11 [Uncaptioned image] (4,1) 9 [Uncaptioned image] (4,2) 2
[Uncaptioned image] (4,2) 2 [Uncaptioned image] (4,2) 2 [Uncaptioned image] (4,2) 4 [Uncaptioned image] (4,2) 4
[Uncaptioned image] (4,2) 3 [Uncaptioned image] (4,2) 1 [Uncaptioned image] (4,2) 1 [Uncaptioned image] (4,2) 1
[Uncaptioned image] (4,3) 3 [Uncaptioned image] (4,3) 3 [Uncaptioned image] (4,3) 2 [Uncaptioned image] (4,4) 1
[Uncaptioned image] (5,0) 1 [Uncaptioned image] (5,0) 1 [Uncaptioned image] (5,1) 4 [Uncaptioned image] (5,1) 4
[Uncaptioned image] (5,2) 3 [Uncaptioned image] (5,2) 1 [Uncaptioned image] (5,2) 3 [Uncaptioned image] (5,2) 1
[Uncaptioned image] (5,3) 1 [Uncaptioned image] (5,3) 2 [Uncaptioned image] (5,3) 2 [Uncaptioned image] (5,3) 1
[Uncaptioned image] (6,0) 1 [Uncaptioned image] (6,1) 5 [Uncaptioned image] (6,2) 1 [Uncaptioned image] (6,2) 1
[Uncaptioned image] (6,2) 3 [Uncaptioned image] (6,2) 1 [Uncaptioned image] (6,3) 2 [Uncaptioned image] (6,3) 2
[Uncaptioned image] (6,3) 2 [Uncaptioned image] (6,3) 1 [Uncaptioned image] (6,4) 1
For the square with two adjacent vertices as a subset, there are actually two different isometric embeddings into $V$, see Remark 2; we indicated their difference by drawing the roof on the square as it appears in the dodecahedron.
3. Convex bodies
We will now refine the construction and show that $S,T$ can be used to construct three-dimensional polytopes $K_{S}$ and $K_{T}$ that are not isometric, but their distribution of isometry classes of planar intersections is identitical. Let us first explain more precisely, what that means: We denote by $\mathcal{P}$ the space of compact planar convex sets modulo Euclidean isometries. For each closed convex set $K\subset\mathbb{R}^{3}$, we consider the natural map
$\Phi_{K}\colon\mathcal{A}(3,2)\longrightarrow\mathcal{P},\qquad H\longmapsto[K\cap H]_{\mathrm{iso}},$
where $\mathcal{A}(3,2)$ is the affine Grassmannian of planes in $\mathbb{R}^{3}$. Note that $\mathcal{A}(3,2)$ carries a standard infinite rotation- and translation-invariant measure $\lambda$ and we call the Borel measure $(\Phi_{K})_{*}(\lambda)$ on $\mathcal{P}$ the distribution of isometry classes of planar sections. Our normalization of $\lambda$ is in such a way that $(\Phi_{K})_{*}(\lambda)(\mathcal{P})={\rm vol}(K)$ for all compact convex bodies.
Theorem 4.
There exists three-dimensional polytopes $K,L\subseteq\mathbb{R}^{3}$, such that
  1. (1)
    $K$ and $L$ are not congruent, but
  2. (2)
    $(\Phi_{K})_{*}(\lambda)=(\Phi_{L})_{*}(\lambda)$.
Proof.
Let $D\subset\mathbb{R}^{3}$ be a regular dodecahedron with vertex set $V$, and let $S,T\subset V$ be two subsets as in Theorem 3: they are not congruent in $\mathbb{R}^{3}$, but their planar statistics agree, i.e. $\mathrm{PS}(S)=\mathrm{PS}(T)$ as multisets of planar congruence classes in the sense of Definition 1.
Let us now explain how we construct polytopes out of $S$ and $T$. For each $X\subset V$ define a convex polytope $K_{X}$ by truncating all vertices $v\in X$ by small congruent triangular caps. We will show that $K_{S}$ and $K_{T}$ are not congruent in $\mathbb{R}^{3}$, but the distributions of isometry classes of their planar sections with respect to the invariant measure on affine planes coincide.
Fix once and for all a small parameter $\varepsilon>0$. For each vertex $v\in V$ let $P_{v}$ be the unique plane that intersects the three edges incident to $v$ in points of distance $\varepsilon$ from $v$, and denote by $H_{v}^{-}$ the closed half-space bounded by $P_{v}$ that contains the barycenter of $D$. For any subset $X\subset V$ we set
$K_{X}:=D\cap\bigcap_{v\in X}H_{v}^{-}.$
By construction $K_{X}$ is obtained from $D$ by cutting off, at each $v\in X$, a small congruent triangular pyramid. In particular, if $g\in\mathrm{Isom}(\mathbb{R}^{3})$ is an isometry with $g(K_{S})=K_{T}$, then $g$ must map the set of truncated vertices of $K_{S}$ onto that of $K_{T}$, and restricts to an isometry $g\colon D\to D$ with $g(S)=T$. Since $S$ and $T$ are not congruent as subsets of $V$ by Theorem 3, it follows that $K_{S}$ and $K_{T}$ are not congruent.
We now compare the distributions of planar sections. For $H\in\mathcal{A}(3,2)$ set
$m(H):=\bigl|\{v\in V:H\cap B(v,\varepsilon)\neq\emptyset\}\bigr|,$
the number of vertex neighbourhoods met by $H$. For each integer $m\geq 0$ define
$A_{m}:=\{H\in\mathcal{A}(3,2):m(H)=m\},$
so that $\mathcal{A}(3,2)=\sqcup_{m\geq 0}A_{m}.$ Let $p_{m}:=\lambda(A_{m})$ and $\lambda_{m}:=p_{m}^{-1}\,\lambda\!\restriction_{A_{m}}$ for all $m$ with $p_{m}>0$. Then, we have
$\lambda=\sum_{m\geq 0}p_{m}\,\lambda_{m},$
a canonical convex decomposition of $\lambda$, and for each $X\subset V$ we obtain a corresponding convex decomposition
$\mu_{X}=\sum_{m\geq 0}p_{m}\,\mu_{X,m},\qquad\mu_{X,m}:=(\Phi_{K_{X}})_{*}(\lambda_{m}).$
To prove $\mu_{S}=\mu_{T}$ it therefore suffices to show that $\mu_{S,m}=\mu_{T,m}$ for all $m\geq 0$.
The cases $m=0$ and $1$. If $H\in A_{0}$, then $H$ misses all balls $B(v,\varepsilon)$ and in particular all caps. The section $K_{X}\cap H$ then coincides with $D\cap H$; it does not depend on $X$. In particular, we can conclude $\mu_{S,0}=\mu_{T,0}.$ Similarly, if $H\in A_{1}$, then $H$ meets exactly one ball $B(v,\varepsilon)$, say at $v$. The local shape of the section $K_{X}\cap H$ in a neighbourhood of $B(v,\varepsilon)$ is completely determined by the local geometry at $v$ and thus only depends on whether $v\in X$ or not. Thus, since $D$ is vertex-transitive and $|S|=|T|$, we obtain $\mu_{S,1}=\mu_{T,1}$.
The case $m\geq 3$. Suppose $H\in A_{m}$ with $m\geq 3$. By choice of $\varepsilon>0$, this implies that $H$ lies in a small neighbourhood of a unique vertex plane $\Pi\subset\mathbb{R}^{3}$ with
$\Pi\cap V=\{v_{1},\dots,v_{m}\},\qquad m\geq 3,$
and that the vertices $\{v_{1},\dots,v_{m}\}$ are precisely those whose balls $B(v_{i},\varepsilon)$ meet $H$. The local configuration of the section $K_{X}\cap H$ is then, up to isometry, completely determined by the isometry type of the inclusion $\Pi\cap X\subset V$ and the position of $H$ in a sufficiently small neighbourhood of $\Pi$.
The isometry group $D$ acts on the set of such local configurations. For a fixed $X\subset V$ the orbits of this action are parametrized precisely by the planar congruence classes $\Pi]_{X}$ appearing in the planar statistic $\mathrm{PS}(X)$: two vertex planes $\Pi,\Pi^{\prime}$ give rise to the same local configuration type if and only if there is an isometry $g$ such that $g(V)=V$ and $g(\Pi\cap X)=\Pi^{\prime}\cap X$.
Fix a $G$-orbit of local configurations $\mathcal{O}$, i.e. an orbit of a planar inclusion $Z\subset V$ with $|Z|\geq 3$. Consider the set of planes $H$ whose local configuration belongs to $\mathcal{O}$. By invariance of $\lambda$ and congruence of all caps and faces, the restriction of $\lambda$ to this set has the same total mass and the same normalized image measure under $\Phi_{X}$, independent of $X$; it depends only on $\mathcal{O}$. Thus for each such orbit $\mathcal{O}$ there is a probability measure $\nu_{\mathcal{O}}$ on $\mathcal{P}$ describing the contribution of one occurrence of $\mathcal{O}$, and the component $\mu_{X,m}$ can be written as a finite convex combination
$\mu_{X,m}=\sum_{\mathcal{O}}c_{\mathcal{O}}(X)\,\nu_{\mathcal{O}},$
where the sum runs over all orbits $\mathcal{O}$ with $m\geq 3$, and $c_{\mathcal{O}}(X)$ is a nonnegative coefficient proportional to the number of vertex planes whose local configuration with respect to $X$ lies in $\mathcal{O}$. By definition, the multiset of classes $\Pi]_{X}$ in $\mathrm{PS}(X)$ records exactly these multiplicities. Thus the coefficients $c_{\mathcal{O}}(X)$ are completely determined by $\mathrm{PS}(X)$.
Since $\mathrm{PS}(S)=\mathrm{PS}(T)$ as multisets of planar congruence classes, it follows that $c_{\mathcal{O}}(S)=c_{\mathcal{O}}(T)$ for every orbit $\mathcal{O}$, and hence $\mu_{S,m}=\mu_{T,m}$ for all $m\geq 3$.
The cases $m=2$. If $H\in A_{2}$, then $H$ meets exactly two balls $B(v,\varepsilon)$, say at vertices $v,w\in V$. The local shape of $K_{X}\cap H$ is then, up to isometry, determined by the orbit type of the unordered pair $\{v,w\}$ under the isometry group of $D$ together with the membership of $v,w$ in $X$. Again, there are only finitely many such orbit types of unordered pairs in $V$, and for each such type all corresponding pairs have the same local geometry. Each such pair belongs to various vertex planes with three or more vertices, so its contribution is again encoded in the planar statistic $\mathrm{PS}(X)$; in particular, the counts of pairs of each orbit type contained in $X$ are determined by $\mathrm{PS}(X)$. Since $\mathrm{PS}(S)=\mathrm{PS}(T)$, the contributions for $m=2$ agree: $\mu_{S,2}=\mu_{T,2}$.
Combining all cases, we have shown that $\mu_{S,m}=\mu_{T,m}$ for every $m\geq 0$. Since
$\mu_{S}=\sum_{m\geq 0}p_{m}\,\mu_{S,m}\quad\text{and}\quad\mu_{T}=\sum_{m\geq 0}p_{m}\,\mu_{T,m}$
with the same weights $p_{m}$, we conclude that $\mu_{S}=\mu_{T}$. This shows that the distributions of isometry classes of planar sections of $K_{S}$ and $K_{T}$ with respect to $\lambda$ coincide, as claimed. ∎
Acknowledgments
We thank Vadim Alekseev for interesting discussions on this topic.
We used AI-generated scripted checks to classify vertex-plane types and to tally planar statistics, and independently cross-checked key computations. We also acknowledge assistance of a large language model for drafting and refactoring code used to generate figures and tables; all mathematical claims and computational results were verified independently by us.

### v2

@@
 Motivated by the question of whether such random planar sections determine the underlying three-dimensional structure, we investigate a discrete analogue in which the ambient structure is the vertex set of a Platonic solid and the observed data are congruence classes of planar intersections.
-For the regular dodecahedron with vertex set $V$, we define the planar statistic of a subset $X\subseteq V$ of vertices as the distribution of isometry types of inclusions $\Pi\cap X\subseteq V$, and ask whether this statistic determines $X$ up to planar congruence.
-We show that this is not the case: there exist two non-congruent $7$-element subsets with identical planar statistics.
+For the regular dodecahedron with vertex set $V$, we define the planar statistic of a subset $X\subseteq V$ of vertices as the distribution of isometry types of inclusions $\Pi\cap X\subseteq V$, and ask whether this statistic determines $X$ up to isometry.
+We show that this is not the case: there exist two non-isometric $7$-element subsets with identical planar statistics.
 As a consequence, there exists two polytopes in $\mathbb{R}^{3}$, whose distribution of isometry classes of two-dimensional intersections is identical, while the polytopes are not themselves isometric.
@@
 We refer to this distribution as the planar statistic of $X$, see Definition 1.
-Our primary question is whether two subsets with identical planar statistics must be congruent in $\mathbb{R}^{3}$.
+Our primary question is whether two subsets with identical planar statistics must be isometric in $\mathbb{R}^{3}$.
 By analogy to results of Mallows and Clark [4] on intercept distributions in the plane, non-uniqueness is expected.

### v3

@@
 Motivated by the question of whether such random planar sections determine the underlying three-dimensional structure, we investigate a discrete analogue in which the ambient structure is the vertex set of a Platonic solid and the observed data are congruence classes of planar intersections.
-For the regular dodecahedron with vertex set $V$, we define the planar statistic of a subset $X\subseteq V$ of vertices as the distribution of isometry types of inclusions $\Pi\cap X\subseteq V$, and ask whether this statistic determines $X$ up to isometry.
+For the regular dodecahedron with vertex set $V$, we define the planar statistic of a subset $X\subseteq V$ of vertices as the distribution of isometry types of inclusions $\Pi\cap X\subseteq\Pi\cap V\subseteq V$, and ask whether this statistic determines $X$ up to isometry.
 We show that this is not the case: there exist two non-isometric $7$-element subsets with identical planar statistics.
-As a consequence, there exists two polytopes in $\mathbb{R}^{3}$, whose distribution of isometry classes of two-dimensional intersections is identical, while the polytopes are not themselves isometric.
+As a consequence, there exist two polytopes in $\mathbb{R}^{3}$, whose distribution of isometry classes of two-dimensional intersections is identical, while the polytopes are not themselves isometric.
 This result is an analogue of classical non-uniqueness phenomena in geometric tomography.
@@
 Theorem 3 shows that such a non-congruent pair $(S,T)$ exists inside the vertex–plane incidence structure of the dodecahedron.
-In the second step we again encode this discrete data into convex bodies by a local modification along selected faces.
+In the second step we again encode this discrete data into convex bodies by a local modification at selected vertices.
 Starting from $D$, we cut off a small congruent triangular pyramid at each vertex $v\in X$, obtaining a polytope $K_{X}$ for $X\subseteq V$.
 When an affine plane $H$ meets $D$ near some vertices, the local shape of the section $K_{X}\cap H$ in a neighbourhood of those vertices is completely determined, up to isometry, by the configuration $\Pi\cap X$, where $\Pi$ is the nearby vertex–plane through the intersected vertices.
-The distribution of isometry classes of planar sections of $K_{X}$ can therefore be written as a convex combination of contributions coming from the finitely many planar intersection types $\Pi\cap X\subset V$, with coefficients prescribed precisely by the planar statistic $\mathrm{PS}(X)$.
+The distribution of isometry classes of planar sections of $K_{X}$ can therefore be written as a convex combination of contributions coming from the finitely many planar intersection types $\Pi\cap X\subset\Pi\cap V\subseteq V$, with coefficients prescribed precisely by the planar statistic $\mathrm{PS}(X)$.
 From this viewpoint, Theorem 4 is the exact three-dimensional analogue of the Mallows–Clark theorem: the equality $\mathrm{PS}(S)=\mathrm{PS}(T)$ ensures that the section distributions of $K_{S}$ and $K_{T}$ coincide, while the underlying subsets $S$ and $T$ and hence the patterns of truncated vertices of $D$ are not related by any isometry.
@@
 Figure 1. Schlegel diagram with all 20 vertices labeled.
-For $X\subset V$ the planar statistic $\mathrm{PS}(X)$ is the multiset, over all planes $\Pi$ with $|\Pi\cap V|\geq 3$, of planar congruence classes of subset inclusions $\Pi\cap X\subseteq V$, seen as a point sets in $\mathbb{R}^{3}$.
+For $X\subset V$ the planar statistic $\mathrm{PS}(X)$ is the multiset, over all planes $\Pi$ with $|\Pi\cap V|\geq 3$, of planar congruence classes of subset inclusions $\Pi\cap X\subseteq V$, seen as point sets in $\mathbb{R}^{3}$.
 More precisely, let $D\subset\mathbb{R}^{3}$ be the dodecahedron as above with vertex set $V$.
@@
 For a subset $X\subset V$ and a vertex-plane $\Pi\in\mathcal{H}$ we consider the inclusion of finite subsets $\Pi\cap X\subseteq V$.
-Two vertex-planes $\Pi,\Pi^{\prime}\in\mathcal{H}$ are called planarly congruent with respect to $X$ if there exists a Euclidean isometry $g\in\mathrm{Isom}(\mathbb{R}^{3})$ such that $g(V)=V$ and $g(\Pi\cap X)=\Pi^{\prime}\cap X$.
+Two vertex-planes $\Pi,\Pi^{\prime}\in\mathcal{H}$ are called planarly congruent with respect to $X$ if there exists a Euclidean isometry $g\in\mathrm{Isom}(\mathbb{R}^{3})$ such that $g(V)=V$, $g(\pi\cap V)=\Pi^{\prime}\cap V$ and $g(\Pi\cap X)=\Pi^{\prime}\cap X$.
 We denote the corresponding equivalence relation on $\mathcal{H}$ by $\Pi\sim_{X}\Pi^{\prime}$ and write $\Pi]_{X}$ for the equivalence class of $\Pi$.
@@
 The planar statistic of $X\subset V$ is the multiset $\mathrm{PS}(X)$ of planar congruence classes $\Pi]_{X}$ as $\Pi$ ranges over all vertex-planes in $\mathcal{H}$, counted with multiplicity.
-In other words, $\mathrm{PS}(X)$ records, for each vertex-plane $\Pi$, the isometry type in $\mathbb{R}^{3}$ of the inclusion $\Pi\cap X\subseteq V$, and forgets the actual position of $\Pi$.
+In other words, $\mathrm{PS}(X)$ records, for each vertex-plane $\Pi$, the isometry type in $\mathbb{R}^{3}$ of the inclusion $\Pi\cap X\subseteq\Pi\cap V\subseteq V$, and forgets the actual position of $\Pi$.
 Remark 2.
@@
 isometric subsets of $V$ yield the same planar statistics.
-In view of the applications in shape recognition described in the introduction, our initial hope was that the planar statistics $\mathrm{PS}(X)$ allows to recover $X$, but we show that this is not the case.
+In view of the applications in shape recognition described in the introduction, our initial hope was that the planar statistics $\mathrm{PS}(X)$ allows one to recover $X$, but we show that this is not the case.
 Our main result is the following theorem:
@@
 3. Convex bodies
-We will now refine the construction and show that $S,T$ can be used to construct three-dimensional polytopes $K_{S}$ and $K_{T}$ that are not isometric, but their distribution of isometry classes of planar intersections is identitical.
+We will now refine the argument and show that $S,T$ can be used to construct three-dimensional polytopes $K_{S}$ and $K_{T}$ that are not isometric, but their distribution of isometry classes of planar intersections is identitical.
 Let us first explain more precisely, what that means: We denote by $\mathcal{P}$ the space of compact planar convex sets modulo Euclidean isometries.
@@
 Note that $\mathcal{A}(3,2)$ carries a standard infinite rotation- and translation-invariant measure $\lambda$ and we call the Borel measure $(\Phi_{K})_{*}(\lambda)$ on $\mathcal{P}$ the distribution of isometry classes of planar sections.
-Our normalization of $\lambda$ is in such a way that $(\Phi_{K})_{*}(\lambda)(\mathcal{P})={\rm vol}(K)$ for all compact convex bodies.
+Our normalization of $\lambda$ is in such a way that $(\Phi_{K})_{*}(\lambda)(\mathcal{P}\setminus\{\varnothing\})=W_{1}(K)$ for a compact convex set $K$, where $W_{1}$ is the first quermassintegral.
 Theorem 4.
-There exists three-dimensional polytopes $K,L\subseteq\mathbb{R}^{3}$, such that
+There exist three-dimensional polytopes $K,L\subseteq\mathbb{R}^{3}$, such that
   1. (1)

### v4

@@
 Motivated by the question of whether such random planar sections determine the underlying three-dimensional structure, we investigate a discrete analogue in which the ambient structure is the vertex set of a Platonic solid and the observed data are congruence classes of planar intersections.
-For the regular dodecahedron with vertex set $V$, we define the planar statistic of a subset $X\subseteq V$ of vertices as the distribution of isometry types of inclusions $\Pi\cap X\subseteq\Pi\cap V\subseteq V$, and ask whether this statistic determines $X$ up to isometry.
+For the regular dodecahedron with vertex set $V$, we define the planar statistic of a subset $X\subseteq V$ of vertices as the distribution of isometry types of inclusions $\Pi\cap X\subseteq\Pi\cap V\subseteq V$, and ask whether this statistic determines $X\subset V$ up to isometry.
 We show that this is not the case: there exist two non-isometric $7$-element subsets with identical planar statistics.
@@
 High-resolution imaging techniques in biological systems frequently produce planar sections of inherently three-dimensional structures.
-A notable example appears in organoid research, where multicellular aggregates are cultivated in three-dimensional media, yet standard microscopy visualizes them in two-dimensional optical slices, [2, 3, 6].
-When we think of one two-dimensional slice through a three-dimensional space filled with more or less identical objects each in a random position and orientation, then we effectively obtain a distribution of random slices of one fixed object without information on what slice exactly we are looking at.
-More precisely, the resulting observed data consists of a distribution of planar point configurations or shapes up to planar congruence.
-A fundamental question, both theoretically and for interpretation of imaging data, is whether this distribution encodes the underlying three-dimensional structure uniquely.
-The problem is rather different from problems in $X$-ray image analysis or geometric tomography, where the position of the intersecting plane is part of the data that can be analysed.
-The aim of this work is to study this question in a highly controlled discrete setting.
-We fix a finite vertex set $V$ of a Platonic solid in $\mathbb{R}^{3}$ and consider subsets $X\subset V$ as discrete models for three-dimensional structures.
-Planes intersecting $V$ produce subset inclusions $\Pi\cap X\subseteq V$ of planar sets.
-Only the Euclidean congruence types of these planar configurations are assumed observable, and each plane is unlabelled.
-Thus the observable information associated to $X$ is merely the distribution of planar congruence classes of all such intersections.
-We refer to this distribution as the planar statistic of $X$, see Definition 1.
-Our primary question is whether two subsets with identical planar statistics must be isometric in $\mathbb{R}^{3}$.
-By analogy to results of Mallows and Clark [4] on intercept distributions in the plane, non-uniqueness is expected.
-We establish such non-uniqueness in the case of the regular dodecahedron, see Theorem 3.
-This also implies that there exist two three-dimensional convex bodies whose distribution of isometry classes of two-dimensional intersections is identical, while the two convex bodies are not themselves isometric, see Theorem 4.
+A prominent example appears in organoid research, where multicellular aggregates are cultivated in three-dimensional media, yet standard microscopy visualizes them in two-dimensional optical slices, [4, 5, 11].
+In many experimental pipelines, one does not obtain a calibrated full-volume scan, but rather a large collection of optical sections or histological slices taken at (approximately) random offsets and orientations.
+From a mathematical viewpoint, such data amount to a distribution of planar configurations, observed only up to planar congruence.
+This leads to a basic inverse question: does the distribution of random planar sections identify the underlying three-dimensional structure uniquely?
+The problem is conceptually different from classical geometric tomography, where the position of the slicing plane is part of the measured data; here we treat the slice as unlabelled, and only the isometry class of the resulting planar configuration is retained.
+The classical entry point in stereology is Wicksell’s corpuscle problem for spheres [15]. For a single sphere of radius $r$, a plane at offset $u\in[-r,r]$ cuts a disk of radius
+$\rho=\sqrt{r^{2}-u^{2}}\in[0,r].$
+If the plane offset is sampled uniformly, the induced (non-normalized) section-radius measure on $(0,r)$ is
+$\mu_{r}(d\rho)=\frac{2\rho}{\sqrt{r^{2}-\rho^{2}}}\,d\rho,\qquad\mu_{r}([0,r])=2r.$
+Passing from a single sphere to a population, let $f(r)$ be the density of sphere radii in $\mathbb{R}^{3}$ and let $g(\rho)$ be the density of observed section radii.
+Then $g$ is related to $f$ by an Abel-type integral transform [15, 3]:
+$g(\rho)=2\rho\int_{r=\rho}^{\infty}\frac{f(r)}{\sqrt{r^{2}-\rho^{2}}}\,dr,\qquad\rho>0,$
+and one can invert this relation and recover $f$ from $g$ via
+$f(r)=-\frac{1}{\pi}\frac{d}{dr}\int_{\rho=r}^{\infty}\frac{g(\rho)}{\sqrt{\rho^{2}-r^{2}}}\,d\rho,\qquad r>0.$
+In particular, in the spherical case the radius distribution is uniquely determined by the section statistics.
+Beyond spheres, identifiability becomes subtle.
+For particles of the form $sK$ (random orientation and scale $s$), section functionals such as area lead to one-dimensional integral equations, and the scale distribution can often be identified [12, 14].
+Related stereological methods and estimators also exist for polyhedral shapes such as cubes [9, 10].
+On the other hand, there are genuine non-identifiability results: Cruz-Orive, solving a conjecture of Moran, exhibited distinct distributions of ellipsoids with identical induced distributions of unoriented planar sections [1, 8].
+The aim of this work is to study such identifiability questions at first in a highly controlled model, motivated by random-slice data in microscopy.
+We fix a finite vertex set $V$ of a Platonic solid in $\mathbb{R}^{3}$ and regard subsets $X\subset V$ as discrete proxies for three-dimensional microstructures.
+Each vertex-plane $\Pi$ determines an inclusion $\Pi\cap X\subseteq\Pi\cap V\subseteq V$, and the only observable information is the Euclidean congruence type of this inclusion.
+Aggregating over all vertex-planes, we obtain a distribution of planar congruence classes; we call it the planar statistic of $X$, see Definition 1.
+Our guiding question is whether the planar statistic determines $X$ up to isometry in $\mathbb{R}^{3}$.
+In analogy with non-uniqueness phenomena for intercept and chord-length distributions in the plane [6], one expects that low-dimensional section statistics can fail to be injective.
+Our main theorem confirms this expectation in the case of the regular dodecahedron: we exhibit two non-congruent $7$-element subsets $S,T\subset V$ with identical planar statistics, see Theorem 3.
+In particular, even in an idealized, noise-free regime with isotropic plane sampling, uniformly sampled slices need not suffice to distinguish the underlying structure.
+As a geometric consequence, we construct two polytopes in $\mathbb{R}^{3}$ whose distributions of isometry classes of unoriented planar sections coincide, although the bodies themselves are not isometric; see Theorem 4.
+So, while the result of Cruz-Orive [1] shows non-identifiability for distributions of convex sets, our construction yields non-identifiability already for single polytopes.
+This result is an analogue of classical non-uniqueness phenomena in geometric tomography [2, 13].
 Let us explain our approach and the relation to previous work in more detail.
@@
 To each side in $X$, respectively in $Y$, one attaches a small congruent isosceles triangular hat pointing outward, thereby turning the octagon into two convex dodecagons $K_{X}$ and $K_{Y}$.
-The caps are chosen so small that every chord of $K_{X}$ or $K_{Y}$ either coincides with a chord of the original octagon, or differs from it only in a tiny segment inside one of the caps.
 The contribution of a cap to the chord–length distribution depends only on which side it sits on, so the matching of unordered pairs of sides in $X$ and $Y$ forces the chord–length distributions of $K_{X}$ and $K_{Y}$ to coincide, although the polygons themselves are not congruent.
@@
 The first, combinatorial step is the search for two subsets $S,T\subset V$ with identical planar statistics in the sense of Definition 1.
-Equivalently, for every vertex–plane $\Pi$ the inclusion $\Pi\cap S\subset V$ has, up to isometry of $D$, the same type and multiplicity distribution as $\Pi\cap T\subset V$.
+Equivalently, for every vertex–plane $\Pi$ the inclusion $\Pi\cap S\subset\subset\Pi\cap V\subset V$ has, up to isometry of $D$, the same type and multiplicity distribution as $\Pi\cap T\subset\Pi\cap V\subset V$.
 Theorem 3 shows that such a non-congruent pair $(S,T)$ exists inside the vertex–plane incidence structure of the dodecahedron.
@@
 Starting from $D$, we cut off a small congruent triangular pyramid at each vertex $v\in X$, obtaining a polytope $K_{X}$ for $X\subseteq V$.
-When an affine plane $H$ meets $D$ near some vertices, the local shape of the section $K_{X}\cap H$ in a neighbourhood of those vertices is completely determined, up to isometry, by the configuration $\Pi\cap X$, where $\Pi$ is the nearby vertex–plane through the intersected vertices.
+When an affine plane $H$ meets $D$ near some vertices, the local shape of the section $K_{X}\cap H$ in a neighbourhood of those vertices is completely determined, up to isometry, by the configuration $\Pi\cap X\subset\Pi\cap V\subset V$ , where $\Pi$ is the nearby vertex–plane through the intersected vertices.
 The distribution of isometry classes of planar sections of $K_{X}$ can therefore be written as a convex combination of contributions coming from the finitely many planar intersection types $\Pi\cap X\subset\Pi\cap V\subseteq V$, with coefficients prescribed precisely by the planar statistic $\mathrm{PS}(X)$.
-From this viewpoint, Theorem 4 is the exact three-dimensional analogue of the Mallows–Clark theorem: the equality $\mathrm{PS}(S)=\mathrm{PS}(T)$ ensures that the section distributions of $K_{S}$ and $K_{T}$ coincide, while the underlying subsets $S$ and $T$ and hence the patterns of truncated vertices of $D$ are not related by any isometry.
-Beyond the discrete-geometric questions addressed here, our section statistics interface naturally with classical inverse problems in geometric tomography and integral geometry [1, 7], where distributions of slices, chords, and projections play a central role.
-Related stereological ideas go back at least to Wicksell’s corpuscle problem [8] and remain fundamental in modern microscopy when only section data are available.
+From this viewpoint, Theorem 4 is (upon replacing faces by vertices)the exact three-dimensional analogue of the Mallows–Clark theorem: the equality $\mathrm{PS}(S)=\mathrm{PS}(T)$ ensures that the section distributions of $K_{S}$ and $K_{T}$ coincide, while the underlying subsets $S$ and $T$ and hence the patterns of truncated vertices of $D$ are not related by any isometry.
 2. Main results and model
@@
 where $\mathcal{A}(3,2)$ is the affine Grassmannian of planes in $\mathbb{R}^{3}$.
-Note that $\mathcal{A}(3,2)$ carries a standard infinite rotation- and translation-invariant measure $\lambda$ and we call the Borel measure $(\Phi_{K})_{*}(\lambda)$ on $\mathcal{P}$ the distribution of isometry classes of planar sections.
-Our normalization of $\lambda$ is in such a way that $(\Phi_{K})_{*}(\lambda)(\mathcal{P}\setminus\{\varnothing\})=W_{1}(K)$ for a compact convex set $K$, where $W_{1}$ is the first quermassintegral.
+Note that $\mathcal{A}(3,2)$ is naturally a fibre bundle over $\mathbb{R}{\rm P}(2)$ with fibre $\mathbb{R}$, where the fibre over a line $u\in\mathbb{R}{\rm P}(2)$ is the set of all planes with normal $u$, where the element in the fibre determines the offset.
+A natural measure $\lambda$, invariant under isometries of $\mathbb{R}^{3}$, can then be constructed as the product of the surface area measure on $\mathbb{R}{\rm P}(2)$ and the Lebesgue measure on the fibre $\mathbb{R}$.
+With this normalization of $\lambda$, we obtain $(\Phi_{K})_{*}(\lambda)(\mathcal{P}\setminus\{\varnothing\})=6\cdot W_{2}(K)$ for a compact convex set $K$, where $W_{2}$ is the second quermassintegral.
+We call the Borel measure $(\Phi_{K})_{*}(\lambda)$ on $\mathcal{P}$ the distribution of isometry classes of planar sections.
 Theorem 4.
@@
 Acknowledgments
-We thank Vadim Alekseev for interesting discussions on this topic.
+We thank Vadim Alekseev and Leon Renkin for interesting discussions on this topic.
 We used AI-generated scripted checks to classify vertex-plane types and to tally planar statistics, and independently cross-checked key computations.